# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v1

@@ 4.6. Large Legal Corpora
Niklaus et al. (2024) present the MultiLegalPile, the largest open-source multilingual legal corpus available, totaling 689GB and spanning 17 jurisdictions across 24 languages. This extensive dataset is designed to facilitate training of LLMs within the legal domain, featuring diverse legal text types including case law, legislation, and contracts, predominantly in English due to the integration of the ‘Pile of Law’ (Henderson et al., 2022) dataset. Through careful regex-based filtering from the mC4 corpus and manual reviews, the team ensures high precision in legal content selection. The corpus, efficiently compressed using XZ and formatted in JSONL, supports comprehensive NLP research and modeling, emphasizing its broad applicability in advancing legal AI technologies.
@@ 5. Legal Language Models And Methods For Legal Domain Adaptation
In the fast-moving field of NLP, LLMs have become a key tool for processing and understanding large amounts of unstructured text data. These models, initially trained on broad datasets like Wikipedia, have shown great skill across various language tasks. Building on this success, the legal technology community is increasingly interested in using these powerful models for Legal NLP applications. This involves adapting these general-domain models to legal texts and further training them on specialized legal documents. Such efforts aim to reduce the domain gap and customize the models to better understand the complex language used in legal documents. In this section, we will explore how these models are being adapted and applied within the legal domain to enhance Legal NLP applications.
In this section, following the methodology of this survey, we studied all peer-reviewed LMs or methods. However, due to the significant challenges present in the legal domain, there are many legal LMs that have not undergone peer review. Given the scarcity of adequate peer-reviewed resources, our research has focused on the investigation of, in order of priority, the peer-reviewed sources, then the most well-known and widely used non-peer-reviewed legal LMs. Despite their lack of formal peer review, these models have gained considerable attention and usage in the field.
@@ 5.1. Language Models
Chalkidis et al. (2020) present an in-depth analysis of applying BERT (Devlin et al., 2019), a pre-trained language model, in the legal domain, showcasing the need for domain-specific adaptation to enhance performance on legal NLP tasks. They explore three strategies: using standard BERT directly, further pre-training on legal corpora, and pre-training from scratch with legal-specific data. Their study found that both further pre-training and pre-training from scratch generally outperform the use of BERT directly. They introduce legal-bert, a specialized family of models optimized for legal text, which includes versions for varied computational capacities and demonstrates competitive performance with a lower environmental impact.
Xiao et al. (2021) introduce Lawformer, a Longformer-based (Beltagy et al., 2020) language model adapted for Chinese legal texts, designed to handle extensive document lengths common in legal data. Recognizing the limitation of standard PLMs with shorter token capacities, Lawformer employs a unique combination of sliding window, dilated sliding window, and global attention mechanisms to efficiently process long texts, making it suitable for legal AI tasks like judgment prediction and LQA. Pre-trained on a vast corpus of Chinese legal documents segmented into criminal and civil cases, Lawformer integrates complex sequential dependencies across tokens using these attention techniques, enhancing model performance for legal-specific tasks.
In the development of specialized NLP tools for Arabic legal texts, a model specifically tailored to the unique linguistic features of Arabic jurisprudence was designed, introducing AraLegal-BERT (Al-qurishi et al., 2022) midway through this innovation. This model enhances NLP applications within the legal field by adapting BERT (Devlin et al., 2019) technology to Arabic’s specific content needs, involving pre-training BERT from scratch using a broad range of legal documents, including legislative materials and contracts.
@@
Shi et al. (2024) develop Legal-LM, a specialized language model tailored for Chinese legal consulting, enhanced with a KG to address domain-specific challenges such as data veracity and non-expert user interaction. The framework involves several steps: extensive pre-training on a rich corpus of legal texts integrated with a legal KG, keyword extraction and Direct Preference Optimization to refine responses, and the use of an external legal knowledge base for data retrieval and response validation. This multi-faceted approach ensures that Legal-LM not only comprehends complex legal language but also generates precise and user-aligned legal advice.
@@ 5.2. Methods for Improving In-Domain Adaptability of Legal Language Models
Li et al. (2022a) explore a novel adaptation of LMs for the legal domain by integrating domain-specific unsupervised data from public legal forums to optimize prefix domain adaptation, a parameter-efficient learning approach that trains only about 0.1% of the model’s parameters. They introduce a training methodology where a deep prompt is specifically tuned using a domain-adapted prefix from legal forums and then utilized in various legal tasks, demonstrating improved few-shot performance compared to full model tuning methods like legal-bert (Chalkidis et al., 2020). This approach significantly reduces computational overhead while maintaining or exceeding performance metrics across multiple legal tasks, suggesting an efficient and scalable model for legal NLP applications.
Mamakas et al. (2022) explore strategies for adapting pre-trained transformers to cope with the challenges of long legal texts within the LexGLUE benchmark, focusing on extending input capabilities and enhancing efficiency. They modify Longformer (Beltagy et al., 2020), originally extending up to 4,096 sub-words, to process up to 8,192 sub-words by reducing local attention window size and incorporating a global token at the end of each paragraph to facilitate information flow across longer texts. Additionally, they adapt legal-bert (Chalkidis et al., 2020) to employ TF-IDF representations to manage longer documents effectively, introducing variants like TF-IDF-SRT-LegalBERT, which deduplicates and sorts sub-words by TF-IDF scores, and TF-IDF-EMB-LegalBERT, which incorporates a TF-IDF embedding layer. These adaptations aim to combine the robust capabilities of transformers with the practical requirements of handling extensive legal documents, surpassing the performance of traditional linear classifiers while maintaining computational efficiency.
@@ 6. Open Research Challenges
ORC2: Interpretability and Explainability
@@
ORC3: Transparent Annotation
Transparently annotated datasets are rare in the field of legal NLP. Often, studies mention the involvement of expert annotators for tasks such as classification, question answering, or prediction, but fail to provide details about the annotators’ backgrounds or the specific annotation processes used. To develop unbiased legal NLP systems, it is necessary to document the dataset creation process thoroughly. This includes providing detailed descriptions of the annotation procedures and the qualifications of the annotators involved, which is essential for ensuring the reliability and fairness of the systems trained on these datasets. Researchers need to prioritize transparency to build trust and allow for effective evaluation in the legal NLP field.
ORC4: Multilingual Capabilities
@@
In legal NLP, enhancing multilingual capabilities remains an underdeveloped area. While efforts like MultiLegalPile (Niklaus et al., 2024) have begun to address this, there remains a gap in research for many languages, including but not limited to Persian and Arabic. These limitations significantly restricts the application of legal NLP across diverse legal systems worldwide, which is important for broader accessibility. Multilingual capabilities introduce unique challenges for legal NLP models, primarily due to the distinct linguistic structures of each language, which often require extensive fine-tuning to ensure accuracy and relevancy in legal contexts. Furthermore, each legal system possesses its own set of terms and document standards, which can vary dramatically from one language to another. Therefore, expanding research into these and other underserved languages is essential for making NLP tools universally applicable and effective.
ORC5: Ontology
@@
The use of ontologies in the legal domain is relatively sparse, yet it holds considerable potential to enhance the robustness of AI methodologies. Ontology or KG can also enable the AI models to draw accurate inferences regarding the relationship between the terms and thereby better understand and process complicated legal texts. This approach could advance the capability of AI systems to handle complex legal reasoning and decision-making processes. However, utilizing ontology in legal NLP faces unique challenges. The complexity of legal language and the concept of ‘open texture’, where the meaning of legal terms can evolve over time, complicate the creation of static ontological models (Mommers, 2010). Legal ontologies must be dynamic, reflecting changes in law and its interpretation over time. Additionally, the integration of real-world and legal concepts within ontologies presents further complexity, as it requires accommodating both legal terms and their relevant real-world contexts (Mommers, 2010).
ORC6: Pre-processing Legal Text
@@
Pre-processing legal texts is challenging due to the distinct nature of legal documents. Existing legal corpora are often contained of raw texts that require extensive cleaning and transformation to become suitable for ML models. Additionally, legal documents can include complex nested structures, like clauses within clauses, and cross-references to other legal cases, statutes, or provisions, making it difficult to break them into coherent units for analysis. These challenges of legal documents make it impractical to directly fine-tune LMs on these raw datasets without substantial pre-processing. This requirement complicates the use of large legal datasets, making them hard to convert into formats that NLP models can readily process and learn from. Without addressing these specific complexities, fine-tuning LMs on raw legal data becomes impractical, limiting the effectiveness of legal NLP applications.
ORC7: Reinforcement Learning from Human Feedback (RLHF)
@@
The use of RLHF within the legal domain is notably scarce. Currently, there is only one peer-reviewed work (Nguyen et al., 2021) available that explores this approach. This indicates a significant opportunity for research and development in this area, as RLHF could potentially enhance NLP’s capability to learn and make decisions based on complex legal data under human guidance. Further exploration into this method could lead to more responsive and adaptable legal NLP systems. However, due to the complex nature of legal reasoning and the need for accurate legal knowledge in human feedback phase, RLHF’s integration into legal NLP poses some challenges. Therefore, on the human feedback side, legal experts such as lawyers and judges must provide guidance to ensure the AI accurately interprets and applies complex legal concepts.
ORC8: Expanding Legal Domain Coverage
@@
ORC9: Small Language Models (SLMs)
Research into SLMs specific to the legal domain is notably absent. Addressing this gap could lead to more efficient, resource-conscious solutions that still maintain high performance in legal text processing and analysis. The development of SLMs tailored for legal applications could revolutionize the accessibility and scalability of legal NLP tools.
ORC10: Domain-Specific Efficient Fine-Tuning
Domain-specific efficient fine-tuning within the legal field is an underexplored area, with only two known studies (Louis et al., 2024; Li et al., 2022a) addressing it. Legal texts consist of complex structures and specialized words that standard LMs may not capture without significant adaptation. Additionally, the legal domain cover a vast array of document types, such as case law, statutes, and contracts, each requiring tailored approaches for effective model application. This diversity makes it imperative to develop fine-tuning strategies that do not only adapt a model generally but rather tailor it to understand the differences between these document types. The majority of existing approaches involve fine-tuning the entire model, which can be resource-intensive. More focused research could enable fine-tuning of legal LMs using fewer resources, enhancing the efficiency of deploying these models in practice.
ORC11: Legal Logical Reasoning
Complex legal logical reasoning remains a significant challenge in LJP, particularly in predicting prison terms. Current state-of-the-art methods struggle to achieve high accuracy in this area. This highlights a clear need for enhanced approaches that can effectively handle the complex of legal reasoning.
ORC12: Legal Named Entity Recognition
@@
ORC14: Retrieval-Augmented Generation
@@
ORC15: Automated Legal Assistance System
To decrease human error and the costs of legal services, there’s a need for a comprehensive automated legal assistance system. This system should span all tasks within the legal NLP field, from question answering to judgment prediction, and cater to different legal specializations like civil and financial law, across multiple languages from English to Persian. Developing an accurate LLM pre-trained on a vast, diverse dataset free from biases and unfairness is crucial. This ensures that the automated legal services can reliably and equitably address a wide range of legal issues.
@@
Open Research Challenges LQA LJP LTC LDS Legal NER LLMs Corpora
ORC1: Bias and Fairness
ORC2: Interpretability and Explainability
ORC3: Transparent Annotation
ORC4: Multilingual Capabilities
ORC5: Ontology
ORC6: Pre-processing Legal Text
ORC7: RLHF
ORC8: Expanding Legal Domain Coverage
ORC9: SLMs
ORC10: Domain-Specific Efficient Fine-Tuning
ORC11: Legal Logical Reasoning
ORC12: Legal NER
ORC13: Stochastic Parrots
ORC14: RAG
ORC15: Automated Legal Assistance System
@@ 7. Conclusion
Advances in AI and NLP have improved Legal NLP techniques and models. These improvements help better meet the needs of laypersons in legal matters and ease the workload of legal professionals. This survey provides a comprehensive overview of the advancements in NLP techniques used in the legal domain. Additionally, we discussed the unique characteristics of legal documents. We also reviewed existing datasets and LLMs tailored for the legal domain. Legal NER research spans multiple languages and utilizes diverse methods, from rule-based to BERT-based models. LDS has largely focused on extractive and abstractive methods, including TF-IDF and transformer-based models. In LTC, multi-class classification tasks dominate, with deep learning architectures like CNNs and Bi-LSTMs widely used. LJP primarily focuses on Chinese datasets with deep learning approaches like CNNs. LQA often leverages information retrieval techniques such as BM25, with a significant focus on statutory law. Finally, we explored key ORCs, such as the need for domain-specific fine-tuning strategies, addressing bias and fairness in legal datasets, and the importance of interpretability and explainability. Other challenges include the development of more robust pre-processing techniques, handling multilingual capabilities, and integrating ontology-based methods for more accurate legal reasoning.

### v2

@@ 11.2. Large corpora for pre-training
Niklaus et al. (2024) present the MultiLegalPile, the largest open-source multilingual legal corpus available, totalling 689 GB and spanning 17 jurisdictions across 24 languages. This extensive dataset is designed to facilitate training of LLMs within the legal domain, featuring diverse legal text types including case law, legislation, and contracts, predominantly in English due to the integration of the ‘Pile of Law’ (Henderson et al., 2022) dataset. Through careful regex-based filtering from the mC4 corpus and manual reviews, the team ensures high precision in legal content selection.
@@ 12. Legal Language Models and Methods for Legal Domain Adaptation
In the fast-moving field of NLP, LLMs have become a key tool for processing and understanding large amounts of unstructured text data. These models, initially trained on broad datasets, such as Wikipedia, have shown great skill across various language tasks. Building on this success, the legal technology community is increasingly interested in using these powerful models for legal NLP applications. This involves adapting these general-domain models to legal texts and further training them on specialised legal documents. Such efforts aim to reduce the domain gap and customise the models to better understand the complex language used in legal documents. In this section, we will explore how these models are being adapted and applied within the legal domain to enhance legal NLP applications.
In this section, following the methodology of this survey, we studied all peer-reviewed LMs or methods. However, due to the challenges present in the legal domain, there are many legal LMs that have not undergone peer review. Given the scarcity of adequate peer-reviewed resources, our research has focused on the investigation of, in order of priority, the peer-reviewed sources, then the most well-known and widely used non-peer-reviewed legal LMs. Despite their lack of formal peer review, these models have gained considerable attention and usage in the field.
@@ 12.1. Language Models
Chalkidis et al. (2020) present an in-depth analysis of applying BERT, a pre-trained LM, in the legal domain, showcasing the need for domain-specific adaptation to enhance performance on legal NLP tasks. They explore three strategies: using standard BERT directly, further pre-training on legal corpora, and pre-training from scratch with legal-specific data. Their study found that both further pre-training and pre-training from scratch generally outperform the use of BERT directly. They introduce legal-bert, a specialised family of models optimised for legal text, which includes versions for varied computational capacities and demonstrates competitive performance with a lower environmental impact.
Xiao et al. (2021) introduce Lawformer, a Longformer-based (Beltagy et al., 2020) LM adapted for Chinese legal texts, designed to handle extensive document lengths common in legal data. Recognising the limitation of standard PLMs with shorter token capacities, Lawformer employs a unique combination of sliding window, dilated sliding window, and global attention mechanisms to process long texts, making it suitable for legal AI tasks, such as LJP and LQA. Pre-trained on a vast corpus of Chinese legal documents segmented into criminal and civil cases, Lawformer integrates complex sequential dependencies across tokens using these attention techniques, enhancing model performance for legal-specific tasks.
In the development of specialised NLP tools for Arabic legal texts, a model specifically tailored to the unique linguistic features of Arabic jurisprudence was designed, introducing AraLegal-BERT (Al-qurishi et al., 2022) midway through this innovation. This model enhances NLP applications within the legal field by adapting BERT technology to Arabic’s specific content needs, involving pre-training BERT from scratch using a broad range of legal documents, including legislative materials and contracts.
@@
Shi et al. (2024) develop Legal-LM, a specialised LM tailored for Chinese legal consulting, enhanced with a KG to address domain-specific challenges such as data veracity and non-expert user interaction. The framework involves several steps: extensive pre-training on a rich corpus of legal texts integrated with a legal KG, keyword extraction and Direct Preference Optimisation to refine responses, and the use of an external legal knowledge base for data retrieval and response validation. This multi-faceted approach ensures that Legal-LM not only comprehends complex legal language but also generates precise and user-aligned legal advice.
@@ 12.2. Methods for Improving In-Domain Adaptability of Legal Language Models
Li et al. (2022) explore a novel adaptation of LMs for the legal domain by integrating domain-specific unsupervised data from public legal forums to optimise prefix domain adaptation, a parameter-efficient learning approach that trains only about 0.1% of the model’s parameters. They introduce a training methodology where a deep prompt is specifically tuned using a domain-adapted prefix from legal forums and then utilised in various legal tasks, demonstrating improved few-shot performance compared to full model tuning methods, such as legal-bert (Chalkidis et al., 2020). This approach reduces computational overhead while maintaining or exceeding performance metrics across multiple legal tasks, suggesting an efficient and scalable model for legal NLP applications.
Mamakas et al. (2022) explore strategies for adapting pre-trained transformers to cope with the challenges of long legal texts within the LexGLUE benchmark, focusing on extending input capabilities and enhancing efficiency. They modify Longformer (Beltagy et al., 2020), originally extending up to 4,096 sub-words, to process up to 8,192 sub-words by reducing local attention window size and incorporating a global token at the end of each paragraph to facilitate information flow across longer texts. Additionally, they adapt legal-bert to employ TF-IDF representations to manage longer documents, introducing variants, such as TF-IDF-SRT-LegalBERT, which deduplicates and sorts sub-words by TF-IDF scores, and TF-IDF-EMB-LegalBERT, which incorporates a TF-IDF embedding layer. These adaptations aim to combine the robust capabilities of transformers with the practical requirements of handling extensive legal documents, surpassing the performance of traditional linear classifiers while maintaining computational efficiency.
@@ 13. Open Research Challenges
ORC2: Privacy Concern
Privacy concerns in legal NLP are critical, as models often handle highly sensitive documents such as court records (Yin and Habernal, 2022). Beyond basic anonymization, advanced techniques, such as differential privacy add statistical noise during training to ensure that individual data points have minimal influence on the model’s output. Adversarial training further bolsters privacy by simulating potential attacks and guiding the model to suppress identifiable attributes in its representations. Privacy-preserving fine-tuning methods—such as applying differential privacy during fine-tuning, federated learning, or knowledge distillation—allow models to adapt to new legal tasks without exposing sensitive data. However, implementing these techniques requires careful calibration, as excessive privacy constraints may degrade model accuracy. This ongoing challenge underscores the need for continued research into privacy-preserving methods tailored to the unique demands of the legal domain.
ORC3: Interpretability and Explainability
@@
ORC4: Annotation Process and Transparency
Annotation quality and transparency are critical challenges in legal NLP. A growing discussion in the community focuses on descriptive versus prescriptive annotation (Röttger et al., 2022). Descriptive annotation captures the full spectrum of annotator subjectivity, preserving diverse interpretations of legal texts and proving valuable for interpretative tasks such as legal question answering or contract analysis. In contrast, prescriptive annotation enforces a single, consistent standard that is essential for tasks requiring strict adherence to legal norms, such as LJP or statute classification. The choice between these paradigms depends on the specific task, domain context, and intended downstream application. Equally, transparent annotation practices are rarely documented in legal NLP studies, with many reports omitting details about annotators’ backgrounds and the procedures employed. Comprehensive documentation of the annotation process is important for assessing dataset quality, ensuring fairness, and building trust in legal NLP systems. Ultimately, explicit decisions regarding both the annotation paradigm and transparency measures are necessary to produce reliable, suitable datasets.
ORC5: Scarcity of Reliable Annotated data
One of the key ORCs in legal NLP is the scarcity of annotated data, which limits the development of robust models due to the high cost and expertise required for labelling complex legal texts. Data augmentation can offer a promising approach to mitigate this issue, yet it introduces its own set of challenges. The complexity of legal texts complicates the application of general augmentation techniques. Ensuring the quality of augmented data remains critical, with efforts focusing on preserving semantic and syntactic integrity to prevent models from learning incorrect patterns (Perçin et al., 2022). Additionally, the effectiveness of augmentation varies across models, necessitating tailored strategies, while scalability concerns arise with computationally intensive methods (Ghosh et al., 2023). The predominant focus on English limits applicability to multilingual legal contexts, and in extremely low-resource scenarios, there’s a risk of overfitting to augmented data (Ghosh et al., 2023). Future research should prioritise developing domain-specific augmentation techniques that maintain legal text nuances, explore flexible approaches, enhance scalability, and extend methods to support diverse languages, thereby addressing these persistent challenges in legal NLP.
ORC6: Multilingual Capabilities
@@
In legal NLP, enhancing multilingual capabilities remains an underdeveloped area. While efforts, such as MultiLegalPile (Niklaus et al., 2024) have begun to address this, there remains a gap in research for many languages, including but not limited to Persian and Arabic. These limitations restricts the application of legal NLP across diverse legal systems worldwide, which is important for broader accessibility. Multilingual capabilities introduce unique challenges for legal NLP models, primarily due to the distinct linguistic structures of each language, which often require extensive fine-tuning to ensure accuracy and relevancy in legal contexts. Furthermore, each legal system possesses its own set of terms and document standards, which can vary dramatically from one language to another. Therefore, expanding research into these and other underserved languages is essential for making NLP tools universally applicable.
ORC7: Ontology
@@
The use of ontologies in the legal domain is relatively sparse, yet it holds considerable potential to enhance the robustness of AI methodologies. Ontology or KG can also enable the AI models to draw accurate inferences regarding the relationship between the terms and thereby better understand and process complicated legal texts. This approach could advance the capability of AI systems to handle complex legal reasoning and decision-making processes. However, utilising ontology in legal NLP faces unique challenges. The complexity of legal language and the concept of ‘open texture’, where the meaning of legal terms can evolve over time, complicate the creation of static ontological models (Mommers, 2010). Legal ontologies must be dynamic, reflecting changes in law and its interpretation over time. Additionally, the integration of real-world and legal concepts within ontologies presents further complexity, as it requires accommodating both legal terms and their relevant real-world contexts (Mommers, 2010).
ORC8: Pre-processing Legal Text
@@
Pre-processing legal texts is challenging due to the distinct nature of legal documents. Existing legal corpora are often contained of raw texts that require extensive cleaning and transformation to become suitable for ML models. Additionally, legal documents can include complex nested structures, such as clauses within clauses, and cross-references to other legal cases, statutes, or provisions, making it difficult to break them into coherent units for analysis. These challenges of legal documents make it impractical to directly fine-tune LMs on these raw datasets without substantial pre-processing. This requirement complicates the use of large legal datasets, making them hard to convert into formats that NLP models can readily process and learn from. Without addressing these specific complexities, fine-tuning LMs on raw legal data becomes impractical, limiting the performance of legal NLP applications.
ORC9: Reinforcement Learning from Human Feedback (RLHF)
@@
The use of RLHF within the legal domain is notably scarce. Currently, there is only one peer-reviewed work (Nguyen et al., 2021) available that explores this approach. This indicates an opportunity for research and development in this area, as RLHF could potentially enhance NLP’s capability to learn and make decisions based on complex legal data under human guidance. Further exploration into this method could lead to more responsive and adaptable legal NLP systems. However, due to the complex nature of legal reasoning and the need for accurate legal knowledge in human feedback phase, RLHF’s integration into legal NLP poses some challenges. Therefore, on the human feedback side, legal experts such as lawyers and judges must provide guidance to ensure the AI accurately interprets and applies complex legal concepts.
ORC10: Expanding Legal Domain Coverage
@@
ORC11: Small Language Models (SLMs)
Research into SLMs specific to the legal domain is notably absent. Addressing this gap could lead to more efficient, resource-conscious solutions that still maintain high performance in legal text processing and analysis. The development of SLMs tailored for legal applications could revolutionise the accessibility and scalability of legal NLP tools.
ORC12: Domain-Specific Efficient Fine-Tuning
Domain-specific efficient fine-tuning within the legal field is an underexplored area, with only two known studies (Louis et al., 2024; Li et al., 2022) addressing it. Legal texts consist of complex structures and specialised words that standard LMs may not capture without significant adaptation. Additionally, the legal domain cover a vast array of document types, such as case law, statutes, and contracts, each requiring tailored approaches for effective model application. This diversity makes it imperative to develop fine-tuning strategies that do not only adapt a model generally but rather tailor it to understand the differences between these document types. The majority of existing approaches involve fine-tuning the entire model, which can be resource-intensive. More focused research could enable fine-tuning of legal LMs using fewer resources, enhancing the efficiency of deploying these models in practice.
ORC13: Legal Logical Reasoning
Complex legal logical reasoning remains a challenge in LJP, particularly in predicting prison terms. Current state-of-the-art methods struggle to achieve high accuracy in this area. This highlights a clear need for improved approaches that better handle legal reasoning.
ORC14: Legal Named Entity Recognition
@@
ORC16: Retrieval-Augmented Generation
@@
Open Research Challenges LQA LJP LTC LDS NER AM LLMs Corpora
Bias and Fairness
Privacy Concern
Interpretability and Explainability
Annotation Process and Transparency
Scarcity of Reliable Annotated data
Multilingual Capabilities
Ontology
Pre-processing Legal Text
RLHF
Expanding Legal Domain Coverage
SLMs
Domain-Specific Efficient Fine-Tuning
Legal Logical Reasoning
Legal NER
Stochastic Parrots
RAG
@@ 14. Conclusion
Advances in AI and NLP have improved legal NLP techniques and models. These improvements help better meet the needs of laypersons in legal matters and ease the workload of legal professionals. This survey provides a comprehensive overview of the advancements in NLP techniques used in the legal domain. Additionally, we discussed the unique characteristics of legal documents. We also reviewed existing datasets and LLMs tailored for the legal domain. Legal NER research spans multiple languages and utilises diverse methods, from rule-based to BERT-based models. LDS has largely focused on extractive and abstractive methods, including TF-IDF and transformer-based models. In LTC, multi-class classification tasks dominate, with deep learning architectures, such as CNNs and Bi-LSTMs widely used. LJP primarily focuses on Chinese datasets with deep learning approaches, such as CNNs. LQA often leverages IR techniques such as BM25, with a significant focus on statutory law. Finally, we explored key ORCs, such as the need for domain-specific fine-tuning strategies, addressing bias and fairness in legal datasets, and the importance of interpretability and explainability. Other challenges include the development of more robust pre-processing techniques, handling multilingual capabilities, and integrating ontology-based methods for more accurate legal reasoning.